# The Serial Scaling Hypothesis

### v2

@@ 1 Introduction
The scaling up of machine learning has driven remarkable progress (Achiam et al., 2023; Hoffmann et al., 2022; Dosovitskiy et al., 2021; Kaplan et al., 2020; Krizhevsky et al., 2012), much of this has come from parallel scaling: hardware shifted from CPUs to massively parallel GPUs, architectures moved from RNNs to highly parallelizable Transformers, and algorithms increasingly exploit parallel compute (Dao, 2024; Chowdhery et al., 2023; Jouppi et al., 2017). But some problems stubbornly resist such advances (Aggarwal & Welleck, 2025; Kazemi et al., 2025; Li et al., 2024; Merrill & Sabharwal, 2023b; Marcus, 2018; Chollet, 2019): for a class of tasks, only scaling serial computation—allowing models to perform more sequential steps—yields further progress. It seems not all scaling is created equal.
The necessity of deep or sequential models for some problems has remained mostly theoretical (Telgarsky, 2016; Merrill & Sabharwal, 2025; Chen et al., 2024a). Only recently has scaling test-time computation become recognized, independently of train-time computation (Snell et al., 2024; OpenAI, 2024; Wu et al., 2025). Yet this dichotomy still overlooks the role of parallel vs. serial computation which is central to this paper. The literature on scaling still commonly report both parameter counts and computation FLOPs each as a single number, treating width (parallel) and depth (serial computation) interchangeably (Kaplan et al., 2020; Hoffmann et al., 2022; OpenAI, 2024; Snell et al., 2024).
@@
  • Explains past successes: The breakthrough of deep learning came from increasing network depth (LeCun et al., 2015; Prince, 2023), and Chain-of-Thought (CoT) improves performance by adding more serial steps (Kojima et al., 2022; Li et al., 2024; Merrill & Sabharwal, 2024).
  • Guides practice: The parallel–serial lens guides both theory and practice. Formal results show bounded serial capacity in Transformers (Merrill & Sabharwal, 2023b) and expanded capacity with CoT (Li et al., 2024; Merrill & Sabharwal, 2024), matching empirical gains (Aggarwal & Welleck, 2025; Muennighoff et al., 2025). Our new analysis extends this picture to diffusion: despite thousands of iterations, diffusion models with a $\mathsf{TC}^{0}$ backbone have only constant computation depth, consistent with observed step-count plateaus (Ma et al., 2025; Ravishankar et al., 2024).
@@ 2.1 Complexity Framework for Inherently Serial Problems
We focus our attention on binary decision problems which take in $N$ input tokens and output either “yes” or “no” as depicted in Figure˜2(A). While seemingly limited, this framework includes all problems with discrete responses, since making a discrete choice out of $2^{n}$ options is equivalent to answering $n$ yes/no questions Figure˜2 (C–E) show how cellular automata, many-body mechanics, and open-ended question answering can be represented in this way.
@@
Intuitively, a problem belongs to $\mathsf{TC}$ if and only if it can be solved by a multilayer perceptron (MLP) with polynomial width and polylog depth, where polylog denotes $\sf{poly}(\log N)$ (Parberry & Schnitger, 1988). A problem belongs to $\mathsf{TC}^{0}$ if and only if it can be solved by a constant-depth MLP. We consider $\mathsf{TC}$ problems as parallel problems, since there are sublinear-depth MLPs that solve them.
@@
A common phenomenon in computational complexity is that one can make a neural network shallower at the cost of exponentially larger width (Valiant, 1983; Hajnal et al., 1993; Sherstov, 2007; Williams, 2014; Oliveira & Santhanam, 2015; Eldan & Shamir, 2016; Liang & Srikant, 2016; Cohen et al., 2016; Telgarsky, 2016; Merrill & Sabharwal, 2025). However, such exponential-sized networks are intractable in both memory and computation. We consider them as inefficient solutions in this paper. This exponential depth–width trade-off underlines the importance of characterizing neural networks depth and width separately, which motivates the Serial Scaling Hypothesis.
@@ 2.2 Real-world problems are likely inherently serial
A natural objection to theory: “Do such theoretical obstructions to parallelism arise in the real world?” We say: Yes—particularly if one works in domains involving the need for physical simulation or reinforcement learning. But even outside these areas, it remains highly likely that one will encounter inherently serial problems. Even apparently simple tasks with a polynomial-time algorithm can be inherent serial (Assumption 2.3).
@@ 2.3 Are our models capable of solving inherently serial problems?
Method Parallel?
Solve serial
problem?
MLPs
Transformers
SSMs and Mamba
@@
Diffusion models
($\sf{TC}^{0}$ backbone)
11footnotemark: 1
Table 1: Parallelizable models are limited to parallel problems. Only non-parallelizable models may solve inherently serial problems.
@@
Most modern architectures are designed for scaling parallel computation. Transformers and state-space models (SSMs) (Gu et al., 2021; Gu & Dao, 2023), though often called “sequence” models, process all available input in parallel. This raises a crucial question: “How efficiently do they solve inherently serial problems?”
@@
Diffusion models cannot solve inherently serial problems. A key contribution of the present paper is a new theorem (Theorem˜F.1) on the limitations of diffusion modeling. Although the step-by-step sampling of a diffusion model appear serial, we prove in Section˜4 that a diffusion model with a $\mathsf{TC}^{0}$ backbone remains in $\mathsf{TC}^{0}$, even with infinitely many sampling steps. Thus, despite their stepwise structure, diffusion models are incapable of solving inherently serial problems.
@@ 2.4 Implications of The Serial Scaling Hypothesis
For machine learning practitioners. Many important real-world problems—such as cellular automata, many-body mechanics, and sequential decision-making—are inherently serial. Without sufficient serial computation, solving them with shallow models requires an exponentially large set of weights, which in turn demands exponentially large datasets to train. Neither is affordable in practice. This mismatch—solving inherently serial problems using shallow or parallel models—helps explain why many ML systems generalize poorly beyond their training distributions (Torralba & Efros, 2011; Zhang et al., 2017; Recht et al., 2019; Liang et al., 2023; Mancoridis et al., 2025; Zhang et al., 2025).
@@
For task & benchmark designers. Recognizing the inevitable cost to inherent seriality, serial problems may be reformulated into coarser or approximate problems to reduce serial depth to acceptable levels while remain practically useful. For RL, truncated value functions cap the effective depth while retaining theoretical guarantees and practical utility (Park et al., 2025; Sutton & Barto, 2018; De Asis et al., 2019). For reasoning, complexity theory shows that coarse decision problems can be tractable even when exact solutions are not—for example, primality is in $\mathsf{P}$ while factoring remains hard (Agrawal et al., 2004).
@@
For hardware designers. Massively parallel computing machinery, especially GPU clusters, enabled past progress in deep learning. Future progress in machine learning and computing in general [^2] Amdahl’s law (Amdahl, 1967; Gustafson, 1988) shows that seriality is a hard upper bound on speedups achievable by parallelism. also depends on progress in high-clockrate, sequential computing machinery, with reduced data movement overhead (Kang et al., 2021; Kaur et al., 2024).
@@ 2.5 Related Works
Our hypothesis is similar to the “parallelism tradeoff” hypothesis—that all parallelizable models, irrespective of design, must necessarily fail to solve inherently serial problems (Merrill & Sabharwal, 2023a; b). Our work builds on classical complexity theory, including the “depth-width tradeoff” (Vishkin & Wigderson, 1985), the “work” vs. “depth” contrast  Blelloch (1996), $\mathsf{P}$-completeness (Greenlaw et al., 1995, Ch. 8), and computational irreducibility (Wolfram, 2002).
@@ 3.1 Cellular Automata
Cellular automata are inherently serial. Rule 110 has been proven Turing-complete (Woods & Neary, 2009), so computing the state of a cell $x_{i}$ at row $N$ requires simulating each step in sequence, without shortcuts (Greenlaw et al., 1995, p. 58). This is not unique to Rule 110: many CA problems are $\sf{P}$-complete (Moore, 1997), i.e., not efficiently parallelizable.
What does this mean? Even systems governed by just 8 simple rules can forbid shortcuts: reducing serial computation often incurs an exponential parallel cost, showing that serial and parallel computation are not interchangeable.
@@ 3.2 Many-Body Mechanics
Consider $N$ particles evolving in $\mathbb{R}^{d}$ under forces and/or hard collisions. Given initial positions and momenta at $t=0$, the goal is to predict the particle positions at some later time $t=T$ in a finite-precision representation. [^3] Finite precision reflects practical scientific computation. With unbounded precision, related prediction problems become $\mathsf{PSPACE}$-hard. This is the instance depicted in Figure˜2(D).
Many-body mechanics is inherently serial. Classical mechanics is expressive enough to emulate arbitrary computation. The billiard-ball computer of Fredkin & Toffoli (1982) recreates any Turing machine using nothing but billiard balls elastically reflecting off each other and walls, and Moore (1990) proposes a similar recreation using smooth particle motion in a potential field in $\mathbb{R}^{3}$. Since the problem of simulating general Turing machines is inherently serial (Greenlaw et al., 1995, p. 58), an algorithm that accurately simulates general physical systems can only be done step by step.
@@ 3.3 $\mathsf{P}$-complete Problems
Such problems have been formalized as $\mathsf{P}$-complete problems, of which the most prototypical one is the Circuit Value Problem (CVP) (Ladner, 1975): Given a Boolean circuit with a single output bit, specified as a graph of logic gates their connections, and an input binary string, what is the output bit? The CVP is robust in that many variations remain $\mathsf{P}$-complete (Greenlaw et al., 1995).
@@
This gives us an intuitive rule: If the problem appears to have a linear computational graph similar to the CVP, then it is likely inherently serial. Indeed, this intuition guided us in proving the theorem in Section˜3.4. Furthermore, an extensive list of problems are shown to be $\mathsf{P}$-complete (Greenlaw et al., 1995), thus suggesting that inherently serial problems are common in the wild.
@@ 3.4 Sequential Decision Problems
In policy gradient, the parameters of a policy model are improved via gradient descent on return estimates, which must be unbiased for convergence to the optimal policy [^4] Convergence to suboptimal policies is still possible with biased return estimates (Mu & Klabjan, 2024; Tian et al., 2023).. Here, we show how an unbiased return estimation is inherently serial, implying that with parallel estimation the optimal policy is not guaranteed.
(a) MCTS in Hex board game: Performance improves with more MTCS expansion nodes across all training regimes. Perfect play is only possible with test-time MCTS. Data from Villalobos & Atkinson (2023).
@@
If parallel return estimation is possible, one must be able to access state $i$ (and thus reward $r_{i}$) in fewer than a linear function of $i$ steps. This is unlikely in real-world MDPs, since even simple CA transition rules (Section˜3.1) or physical interactions (Section˜3.2) prevent such shortcuts [^5] As usual, there is an exponential tradeoff available. One can randomly generate a trajectory and check its consistency with the MDP and policy in parallel, but the expected number of trajectories needed before a consistent one is found grows exponentially with length $n$.. While scaling parallel computation via aggregating multiple trajectories can accelerate convergence by reducing variance (Sutton & Barto, 2018, p. 93), accurate return estimation still requires serial computation.
This inherent seriality motivates model-based reinforcement learning, where returns are computed by unrolling an internal model step by step. For instance, Monte Carlo Tree Search (MCTS)—which increases serial computation via tree expansion and reduces return-estimation bias—has achieved superhuman-level performance in diverse board games (Silver et al., 2016; 2017; Schrittwieser et al., 2020). Villalobos & Atkinson (2023) demonstrate consistent improvements in Hex from increasing MCTS expansion nodes (see Figure˜6(a)). Even in model-free RL, adding serial computation with deeper networks significantly outperform wider (more parallel) ones on locomotion tasks Kevin et al. (2025) (see Figure˜6(b)).
@@ 3.5 Reasoning Question Answering
Figure 7: Average benchmark scores over 4 math benchmarks for longer reasoning chains (in different colors) vs. majority voting (as dots along each line). Data from Aggarwal & Welleck (2025).
@@
In mathematics QA, as shown in Figure˜7, Aggarwal & Welleck (2025) demonstrate that sequential scaling with longer reasoning chains consistently outperforms parallel scaling via majority voting controlled for the same token budget. This pattern holds across mathematical datasets of varying difficulty, including AMC, MATH, AIME, and Olympiad-bench.

### v3

@@ 1 Introduction
The scaling up of machine learning has driven remarkable progress (Achiam et al., 2023; Hoffmann et al., 2022; Dosovitskiy et al., 2021; Kaplan et al., 2020; Krizhevsky et al., 2012), much of this has come from parallel scaling: hardware shifted from CPUs to massively parallel GPUs, architectures moved from RNNs to highly parallelizable Transformers, and algorithms increasingly exploit parallel compute (Dao, 2024; Chowdhery et al., 2023; Jouppi et al., 2017). But some problems stubbornly resist such advances (Aggarwal and Welleck, 2025; Kazemi et al., 2025; Li et al., 2024; Merrill and Sabharwal, 2023b; Marcus, 2018; Chollet, 2019): for a class of tasks, only scaling serial computation—allowing models to perform more sequential steps—yields further progress. It seems not all scaling is created equal.
The necessity of deep or sequential models for some problems has remained mostly theoretical (Telgarsky, 2016; Merrill and Sabharwal, 2025; Chen et al., 2024a). Only recently has scaling test-time computation become recognized, independently of train-time computation (Snell et al., 2024; OpenAI, 2024; Wu et al., 2025). Yet this dichotomy still overlooks the role of parallel vs. serial computation, which is central to this paper. The literature on scaling still commonly reports both parameter counts and computation FLOPs each as a single number, treating width (parallel) and depth (serial computation) interchangeably (Kaplan et al., 2020; Hoffmann et al., 2022; OpenAI, 2024; Snell et al., 2024).
@@
  • Explains past successes: The breakthrough of deep learning came from increasing network depth (LeCun et al., 2015; Prince, 2023), and Chain-of-Thought (CoT) improves performance by adding more serial steps (Kojima et al., 2022; Li et al., 2024; Merrill and Sabharwal, 2024).
  • Connects to practice: The parallel–serial lens connects well to practice. Formal results show bounded serial capacity in Transformers (Merrill and Sabharwal, 2023b) and expanded capacity with CoT (Li et al., 2024; Merrill and Sabharwal, 2024), matching empirical gains (Aggarwal and Welleck, 2025; Muennighoff et al., 2025). Our new analysis extends this picture to diffusion: despite thousands of iterations, diffusion models with a $\mathsf{TC}^{0}$ backbone have only constant computation depth, consistent with observed step-count plateaus (Ma et al., 2025; Ravishankar et al., 2024).
@@ 2.1 Complexity Framework for Inherently Serial Problems
We focus our attention on binary decision problems which take in $N$ input tokens and output either “yes” or “no” as depicted in Figure˜2(A). While seemingly limited, this framework includes all problems with discrete responses, since making a discrete choice out of $2^{n}$ options is equivalent to answering $n$ yes/no questions. Figure˜2 (C–E) show how cellular automata, many-body mechanics, and open-ended question answering can be represented in this way.
@@
Intuitively, a problem belongs to $\mathsf{TC}$ if and only if it can be solved by a multilayer perceptron (MLP) with polynomial width and polylog depth, where polylog denotes $\sf{poly}(\log N)$ (Parberry and Schnitger, 1988). A problem belongs to $\mathsf{TC}^{0}$ if and only if it can be solved by a constant-depth MLP. We consider $\mathsf{TC}$ problems as parallel problems, since there are sublinear-depth MLPs that solve them.
@@
A common phenomenon in computational complexity is that one can make a neural network shallower at the cost of exponentially larger width (Valiant, 1983; Hajnal et al., 1993; Sherstov, 2007; Williams, 2014; Oliveira and Santhanam, 2015; Eldan and Shamir, 2016; Liang and Srikant, 2016; Cohen et al., 2016; Telgarsky, 2016; Merrill and Sabharwal, 2025). However, such exponential-sized networks are intractable in both memory and computation. We consider them as inefficient solutions in this paper. This exponential depth–width trade-off underlines the importance of characterizing neural networks depth and width separately, which motivates the Serial Scaling Hypothesis.
@@ 2.2 Real-world problems are likely inherently serial
A natural objection to theory: “Do such theoretical obstructions to parallelism arise in the real world?” We say: Yes—particularly if one works in domains involving the need for physical simulation or reinforcement learning. But even outside these areas, it remains highly likely that one will encounter inherently serial problems. Even apparently simple tasks with a polynomial-time algorithm can be inherently serial (Assumption 2.3).
@@ 2.3 Are our models capable of solving inherently serial problems?
Method Parallel? Solve serial problem?
FF MLPs
FF Transformers
FF SSMs (Mamba)
@@
Diffusion models ($\sf{TC}^{0}$ backbone) 11footnotemark: 1
Table 1: Parallelizable models are limited to parallel problems. Only non-parallelizable models may solve inherently serial problems. $F$ stands for Feedforward.
@@
Most modern architectures are designed for scaling parallel computation. Transformers and state-space models (SSMs) (Gu et al., 2021; Gu and Dao, 2023), though often called “sequence” models, process all available input in parallel. This raises a crucial question: “How efficiently do they solve inherently serial problems?”
@@
Diffusion models cannot solve inherently serial problems. A key contribution of the present paper is a new theorem (Theorem˜F.2) on the limitations of diffusion modeling. Although the step-by-step sampling of a diffusion model appear serial, we prove in Section˜4 that a diffusion model with a $\mathsf{TC}^{0}$ backbone remains in $\mathsf{TC}^{0}$, even with infinitely many sampling steps. Thus, despite their stepwise structure, diffusion models are incapable of solving inherently serial problems.
@@ 2.4 Implications of The Serial Scaling Hypothesis
For machine learning practitioners. Many important real-world problems—such as cellular automata, many-body mechanics, and sequential decision-making—are inherently serial. Without sufficient serial computation, solving them with shallow models requires an exponentially large set of weights, which in turn demands exponentially large datasets to train. Neither is affordable in practice. This mismatch—solving inherently serial problems using shallow or parallel models—helps explain why many ML systems generalize poorly beyond their training distributions (Torralba and Efros, 2011; Zhang et al., 2017; Recht et al., 2019; Liang et al., 2023; Mancoridis et al., 2025; Zhang et al., 2025).
@@
For task & benchmark designers. Recognizing the inevitable cost to inherent seriality, serial problems may be reformulated into coarser or approximate problems to reduce serial depth to acceptable levels while remain practically useful. For RL, truncated value functions cap the effective depth while retaining theoretical guarantees and practical utility (Park et al., 2025; Sutton and Barto, 2018; De Asis et al., 2019). For reasoning, complexity theory shows that coarse decision problems can be tractable even when exact solutions are not—for example, primality is in $\mathsf{P}$ while factoring remains hard (Agrawal et al., 2004).
@@
For hardware designers. Massively parallel computing machinery, especially GPU clusters, enabled past progress in deep learning. Future progress in machine learning and computing in general [^2] Amdahl’s law (Amdahl, 1967; Gustafson, 1988) shows that seriality is a hard upper bound on speedups achievable by parallelism. also depends on progress in high-clockrate, sequential computing machinery, with reduced data movement overhead (Kang et al., 2021; Kaur et al., 2024). A concrete example is wafer-scale hardware (e.g., Cerebras CS-3), which prioritizes extremely high on-chip memory bandwidth (Wang, 2025).
@@ 2.5 Related Works
Our hypothesis is similar to the “parallelism tradeoff” hypothesis—that all parallelizable models, irrespective of design, must necessarily fail to solve inherently serial problems (Merrill and Sabharwal, 2023a; b). Our work builds on classical complexity theory, including the “depth-width tradeoff” (Vishkin and Wigderson, 1985), the “work” vs. “depth” contrast  Blelloch (1996), $\mathsf{P}$-completeness (Greenlaw et al., 1995, Ch. 8), and computational irreducibility (Wolfram, 2002).
@@ 3.1 Cellular Automata
Cellular automata are inherently serial. Rule 110 has been proven Turing-complete (Woods and Neary, 2009), so computing the state of a cell $x_{i}$ at row $N$ requires simulating each step in sequence, without shortcuts (Greenlaw et al., 1995, p. 58). This is not unique to Rule 110: many CA problems are $\sf{P}$-complete (Moore, 1997), i.e., not efficiently parallelizable.
What does this mean? Even systems governed by just 8 simple rules can forbid shortcuts.
@@ 3.2 Many-Body Mechanics
Consider $N$ particles evolving in $\mathbb{R}^{d}$ under forces and/or hard collisions. Given initial positions and momenta at $t=0$, the goal is to predict the particle positions at some later time $t=T$ in a finite-precision representation, which reflects practical scientific computation; with unbounded precision, the prediction problem becomes $\mathsf{PSPACE}$-hard. This is the instance depicted in Figure˜2(D).
Many-body mechanics is inherently serial. Classical mechanics is expressive enough to emulate arbitrary computation. The billiard-ball computer of Fredkin and Toffoli (1982) recreates any Turing machine using nothing but billiard balls elastically reflecting off each other and walls, and Moore (1990) proposes a similar recreation using smooth particle motion in a potential field in $\mathbb{R}^{3}$. Since the problem of simulating general Turing machines is inherently serial (Greenlaw et al., 1995, p. 58), an algorithm that accurately simulates general physical systems can only be done step by step.
@@ 3.3 $\mathsf{P}$-complete Problems
Such problems have been formalized as $\mathsf{P}$-complete problems, of which the most prototypical one is the Circuit Value Problem (CVP) (Ladner, 1975): Given a Boolean circuit with a single output bit, specified as a graph of logic gates and their connections, and an input binary string, what is the output bit? The CVP is robust in that many variations remain $\mathsf{P}$-complete (Greenlaw et al., 1995).
@@
This gives us an intuitive rule: If the problem appears to have a linear computational graph similar to the CVP, then it is likely inherently serial. Indeed, this intuition guided us in proving the theorem in Section˜3.4. Furthermore, an extensive list of problems is shown to be $\mathsf{P}$-complete (Greenlaw et al., 1995), thus suggesting that inherently serial problems are common in the wild.
@@ 3.4 Sequential Decision Problems
In policy gradient, the parameters of a policy model are improved via gradient descent on return estimates, which must be unbiased for convergence to the optimal policy [^3] Convergence to suboptimal policies is still possible with biased return estimates (Mu and Klabjan, 2024; Tian et al., 2023).. Here, we show how an unbiased return estimation is inherently serial, implying that with parallel estimation the optimal policy is not guaranteed.
(a) MCTS in Hex board game: Performance improves with more MTCS expansion nodes across all training regimes. Perfect play is only possible with test-time MCTS. Data from Villalobos and Atkinson (2023).
@@
If parallel return estimation is possible, one must be able to access state $i$ (and thus reward $r_{i}$) in fewer than a linear function of $i$ steps. This is unlikely in real-world MDPs, since even simple CA transition rules (Section˜3.1) or physical interactions (Section˜3.2) prevent such shortcuts [^4] As usual, there is an exponential tradeoff available. One can randomly generate a trajectory and check its consistency with the MDP and policy in parallel, but the expected number of trajectories needed before a consistent one is found grows exponentially with length $n$.. While scaling parallel computation via aggregating multiple trajectories can accelerate convergence by reducing variance (Sutton and Barto, 2018, p. 93), accurate return estimation still requires serial computation.
This inherent seriality motivates model-based reinforcement learning, where returns are computed by unrolling an internal model step by step. For instance, Monte Carlo Tree Search (MCTS)—which increases serial computation via tree expansion and reduces return-estimation bias—has achieved superhuman-level performance in diverse board games (Silver et al., 2016; 2017; Schrittwieser et al., 2020). Villalobos and Atkinson (2023) demonstrate consistent improvements in Hex from increasing MCTS expansion nodes (see Figure˜6(a)). Even in model-free RL, adding serial computation with deeper networks significantly outperform wider (more parallel) ones on locomotion tasks Kevin et al. (2025) (see Figure˜6(b)).
@@ 3.5 Reasoning Question Answering
Figure 7: Average benchmark scores over 4 math benchmarks for longer reasoning chains (in different colors) vs. majority voting (as dots along each line). Data from Aggarwal and Welleck (2025).
@@
In mathematics QA, as shown in Figure˜7, Aggarwal and Welleck (2025) demonstrate that sequential scaling with longer reasoning chains consistently outperforms parallel scaling via majority voting controlled for the same token budget. This pattern holds across mathematical datasets of varying difficulty, including AMC, MATH, AIME, and Olympiad-bench.